# Measuring Information Distortion in Hierarchical Ultra long Novel Generation: The Optimal Expansion Ratio

### v1

@@ Abstract
Writing novels with Large Language Models (LLMs) raises a key question: how much human-authored guidance is needed to generate high-quality, million-word narratives? Existing frameworks (e.g., DOME, Plan&Write, Long Writer) target long texts (10k–100k words), leaving ultra-long generation underexplored.
Building on recent text compression methods (LLMZip, LLM2Vec), we conduct an information-theoretic analysis to quantify semantic distortion under different compression–expansion ratios. We propose a hierarchical two-stage pipeline (outline $\rightarrow$ section outline $\rightarrow$ manuscript) and examine how outline length affects information preservation.
Experiments on ultra long novels show that our mixed two-stage method significantly reduces semantic distortion compared to single-stage generation.
{CJK}
UTF8gbsn
Measuring Information Distortion in Hierarchical Ultra long Novel Generation:
@@ 1 Introduction
We observe a fundamental phenomenon in large-scale novel generation (Li et al., 2022; Su et al., 2022): when a full-length novel of roughly 1 million words is summarized into a 1,000-word outline by an LLM such as ChatGPT and subsequently expanded back into a novel, the result exhibits substantial distortion, loss of detail, and semantic differences. The reconstructed text diverges markedly from the original. By contrast, compressing a 100k-word novel into a 10k-word outline and then expanding it yields less distortion and more faithfully preserves the original semantics.
Existing studies, such as LongWriter, agree that producing novels longer than 100k words with large language models (LLMs) typically relies on an outline-to-novel workflow. The amount of narrative detail an ultra-long novel produced by an LLM can contain is governed by two prompt-side variables: (i) outline length—the total number of prompt tokens, and (ii) detail density—the average number of details per token (e.g., distinct characters, plots, and scenes).
In practice, however, there is a trade-off: An outline prompt with few details and tokens tends to yield a low-quality novel, whereas an outline prompt with extensive details and many tokens improves novel quality. In the latter case, the human author does most of the work.
Our research question is: What is the minimal outline length required to faithfully reconstruct a million-word novel using state-of-the-art LLMs?
Our ultimate goal is to keep the input outline short while still generating a detail-rich novel.
@@ 1.1 Motivation
Ultra-long novels (over 1 million words) enjoy significant popularity on platforms such as WuxiaWorld, Fanqie, and Qidian. Millions of readers on these platforms value extended, intricate novels, making the automatic generation of high-quality, long-form novels an important research direction.
Since the release of ChatGPT in 2022, LLMs such as LLaMA (Grattafiori et al., 2024), DeepSeek (DeepSeek-AI et al., 2025), Qwen (Bai et al., 2023), Gemini (Team et al., 2024), and Claude have advanced text generation. The 2025 launches of GPT-4o (OpenAI et al., 2024) and Gemini 2.0-Flash further improved long-context processing—for example, GPT-4o supports substantially larger input and output windows, as shown in Table  1. These developments bring us closer to end-to-end long-form generation (Liu et al., 2024); however, coherently producing a high-quality, million-word novel from a single short prompt remains a challenge (Mikhaylovskiy, 2023).
@@
Although many prior studies have addressed novels up to 100k words, empirical evidence is lacking to show that methods effective at this scale naturally extend to generating novels of 1 million words. Challenges at the million-word scale—maintaining coherence, thematic consistency, and character development—are qualitatively and quantitatively different.
@@
1.2 Framework
We treat novel generation as encoding (compression into outlines) and decoding (expansion back into novels), aiming to balance minimal input length with narrative richness.
There are many benchmarks for measuring the quality and character consistency of a novel (Chhun, ; Chhun et al., 2024). However, to the best of our knowledge, a benchmark measuring how many details a million-word LLM generated novel should include does not exist.
We approach long-text generation from an information-theoretic perspective, modeling it as a rate–distortion problem (Valmeekam et al., 2023; Shin, ; Farquhar et al., 2024; Iyer and Mahajani, 2025). The primary question is determining the minimal compressed outline length that allows an input novel to be approximately reconstructed without exceeding predefined distortion limits. Ideally, such distortion should reflect human perceptual judgments and aesthetic similarity scores (Huang et al., 2024; Doshi and Hauser, 2024).
Lossy compression methods intentionally discard less critical information to significantly reduce data size, ensuring acceptable quality retention for targeted applications (Dubois et al., 2022; Goyal et al., 2008). These techniques are widely applied in image processing, medical imaging, and machine learning, demonstrating notable efficiency gains and improved data-handling capacities (Hu et al., 2022; Jamil et al., 2023; Theis et al., 2017; Koff and Shulman, 2006). Importantly, lossy compression approaches can also enhance lossless methods by exploiting inherent redundancies and generalized patterns within data (Mentzer et al., 2020; Marzen and DeDeo, 2017).
Inspired by earlier work on document-level novel generation (e.g., Yang et al., 2024), we find that the granularity of intermediate representations—specifically, the length and depth of outlines—strongly influences the quality of generated novels.
Through these observations, we rigorously investigate how compression–expansion ratios and outline depth affect the generation of million-word novels with state-of-the-art LLMs.
@@ 1.3 Contribution
  • We establish a quantitative relationship between outline compression ratios and distortion (detail loss) through experiments. $R=0.01$ is the optimal compression-expansion ratio under our configuration and experiment.
  • We propose and empirically verify that mixed two-stage hierarchical outline decreases distortion in generated novels compared to traditional one-stage approaches.
@@ 1.4 Related Work
Long-text generation is a broad research area that encompasses various text types including articles, reports, code, and stories. Long-form text generation faces several challenges compared to short text generation, including capturing inter-sentence relations, maintaining global discourse structures, ensuring generation coherence, and providing controllability (Que et al., 2024; Kumar et al., 2024; Wu et al., 2025a, b). For general long-text generation, researchers have developed techniques such as retrieval-augmented generation (Lewis et al., 2021), iterative self-refinement (Madaan et al., 2023), and self-lengthening approaches (Quan et al., 2024). Recent advancements in large language models with extended context windows have enhanced capabilities for handling long contexts, but significant challenges remain (Minaee et al., 2025).
Long-novel generation.
Long-novel generation is a specific subfield that focuses on generating coherent novels with consistent plots, characters, and settings (Xie et al., 2023; Mikhaylovskiy, 2023). Key methods include: (1) hierarchical outline-to-draft generation (Wang et al., 2023; Yang et al., 2023; Wang et al., 2024); (2) plan-based approaches (Yao et al., 2019; Guan et al., 2021; You et al., 2023); (3) memory-enhanced techniques (Wang et al., 2024); and (4) AgentWrite (Bai et al., 2024). While these approaches have demonstrated promising results for novels of 10k-100k words, there is currently no empirical evidence that these methods can scale effectively to ultra-long novels of 1 million words or more (Chhun, ; Doshi and Hauser, 2024).
@@
Complementary research treats the LLM itself as an adaptive entropy coder (Shin, ). LLMZip (Valmeekam et al., 2023) and related systems such as DeepZip (Goyal et al., 2018)achieve near-Shannon lossless compression by arithmetic coding over the token probabilities $p(w_{t}\mid w_{<t})$. We focus instead on lossy yet semantically faithful compression: given an outline that omits detail, can an LLM regenerate an engaging 1 m-word novel? Our work therefore analyzes where an outline’s per-token entropy $H_{\text{outline}}$ can lie—above the lossless bound, yet low enough to conserve human effort—while still permitting faithful restoration.
@@
We want to evaluate two dimensions: (i) distortion between original ultra-long novel and generated novel, and (ii) the quality of generated novel. Current benchmarks for comparing distortion between two texts include: traditional algorithms such as Cosine distance, Hamming distance, Jaccard similarity, Levenshtein distance, TF-IDF (Gahman and Elangovan, 2023a; Gomaa and Fahmy, 2013; Singh and Singh, 2020; Wang and Dong, 2020; Gahman and Elangovan, 2023b) and pre-trained task-specific models such as BERTScore, BLEU, ROUGE, METEOR (Guan et al., 2022; Xiao et al., 2024; Yu et al., 2024; Yuan et al., 2023). Current benchmarks for evaluating the quality of generated novels include: aesthetic score, coherence, creativity (Ichmoukhamedov et al., 2024; Venkatraman et al., 2024), consistency, and fluency. Some researchers propose that large language models themselves can also be used as judges to grade generated content (Gu et al., 2025).
@@ 2 Methodology
This work instantiates the general framework with $K=2$, yielding a level-1 “global” outline $X_{1}$ and a level-2 “section” outline $X_{2}$, followed by two mirrored expansion steps $\psi_{2},\psi_{1}$. In this case, we do not use the sum $\sum_{k=0}^{2}H(X_{k})$ to compute rate–distortion R because our ultimate goal is to generate ultra-long novels from merely global outlines; short section outlines and long global outlines are not expected.
We also test the baseline case with depth $K=1$, which means we summarise a global outline directly from the novel.
Through traditional compression algorithm like LZ77 algorithm, one can obtain lossless compressed text with $R\approx 0.5$ (Amir et al., 2024) which should be a lower bound.
In our study, Distortion is defined as a composite metric incorporating commonly-used cosine distance, BERTScore, and the Euclidean distance of the number of details, which we define in evaluation metrics.
Mathematically, deriving an analytical solution to the ratio-compression problem remains challenging due to the non-linear relationships. Thus we design the following pipeline and experiment.
@@
We adopt a two-stage pipeline:
  1. 1.
    Compression: $X=X_{0}\rightarrow X_{1}$ (varying $\alpha_{1}$), followed by $X_{1}\rightarrow X_{2}$ (varying $\alpha_{2}$).
  2. 2.
    Expansion: $X_{2}\rightarrow\hat{X}_{1}\rightarrow\hat{X_{0}}$ to reconstruct the novel.
  3. 3.
    Direct Expansion: $X_{2}\rightarrow\hat{X_{0}}$ to reconstruct the novel directly.
Both $X_{1}$ and $X_{2}$ may be written in plain language or structured JSON.
2.2 Dataset Configuration
Our dataset consists of two parts: Chinese versions of ultra-long novels from Project Gutenberg, and 40 ultra-long Chinese novels from publicly available sources. These 40 novels from the public internet belong to 4 genres: Fantasy, Urban, Romance, and Historical fiction, and 10 in each genres.
Even though current models in 2025 can process contexts of up to 1M tokens, to ensure accurate information extraction, each 1-million-word novel will be divided and normalized into chapters of approximately 5,000 words each.
Linguistic and expressive variations impact human evaluation processes. We conducted translation experiments demonstrating that different language versions have minimal influence on our study’s assessment.
2.3 Sampling Design
We implement a two-stage design :
  1. 1.
    Stage 1 (between-novel): Stratify the frame into four high-level genres— Urban (U), Romance (R), Fantasy (F) and Historical (H)—then select $n_{h}$ novels in stratum $h$ with probability- proportional-to-size (PPS) where the size variable is total words $L_{i}\approx 1m$. We set the overall testing set size to $n=\mathbf{40}$ and allocate by Neyman optimum allocation(Olayiwola et al., 2013):
    $n_{h}=n\,\frac{N_{h}S_{h}}{\sum_{g}N_{g}S_{g}},$
    where $N_{h}$ and $S_{h}$ are, respectively, the frame count and pilot-estimate standard deviation of the distortion metric in stratum $h$. The first-stage weight is $w_{1,hi}=1/\pi_{hi}$ with $\pi_{hi}=n_{h}L_{i}/\sum_{j\in h}L_{j}$.
  2. 2.
    Stage 2 (within-novel): Treat whole chapters as secondary sampling units. For every selected novel $i$ with $M_{i}$ chapters we draw $m_{i}=\mathbf{8}$ chapters by simple random sampling without replacement (SRSWOR). The second-stage weight is $w_{2,ij}=M_{i}/m_{i}$ (or $1$ if $M_{i}\leq 8$).
The combined design weight is $w_{ij}=w_{1,hi}\,w_{2,ij}$.
2.4 Controlled Variables.
Our configuration space is defined by five independent factors: (i) the novel genre selected for evaluation; (ii) the compression ratio applied at each outline level (global and section); (iii) the outline representation format used to condition the model (plain text versus a structured schema); (iv) the LLM model, which is the backbone model; and (v)the LLM as encoder-decoder setup, which is its generation temperature.
In all ablation runs we modify only one factor at a time while holding the remaining four constant. The content(excludes number of words given to generate in each stage) of prompt itself is not a variable in this paper.
Group ID Method
A Human-written original novel
B Translate $\rightarrow$ Reconstruct
C Direct compression-expansion
D LongWriter
Table 2: Methods for $K=1$ generation baseline.
2.5 Evaluation Metrics
We adopt three automatic evaluation metrics:
Traditional Distance. To measure semantic divergence, we calculate the cosine distance and BERTScore between the embedding representations of the original and generated texts.
LLM-as-Judge. GPT-4o, which is different from the the tool Gemini-2.0-Flash to generate novels, is used to evaluate each generated novel against its human-written original, using semantic similarity, plot similarity, character similarity,background similarity, and style similarity.
Structural Metrics. We compute counts of unique characters, distinct scenes, and distinct items. These are compared against the human-authored originals to estimate structural fidelity.
@@ 2.5 Evaluation Metrics
Table 3: Grouped statistics (mean $\pm$ std) for similarity and structural difference metrics, excluding the baseline (B), to identify best-performing configurations. The compression ratio is computed as $R=\alpha_{1}\times\alpha_{2}$. Bolded values indicate the best average performance for each metric among the tested configurations, but do not imply statistical significance. Formal significance testing results are presented in a separate table.
@@ 3 Experiment
This study investigates the impact of hierarchical compression-expansion strategies on ultra-long novel generation using the Gemini 2.0 Flash model. We aim to answer two core questions: (1) How do different compression ratios affect semantic fidelity in reconstructed texts? (2) Does a two-stage outline hierarchy ($K=2$) outperform a single-stage scheme ($K=1$) in balancing information retention and generation quality?
All generations are performed with Gemini 2.0 Flash (temperature = 0.3), using either plain text or JSON-formatted outlines. Gemini is selected due to its large context window and efficiency in handling high-throughput tasks.
@@ 3.1 Baseline Setup ($K=1$)
We begin with an extreme compression scenario: a single 10,000-word outline is generated from the full novel ($\alpha_{1}\approx 0.01$) and then used as the sole input for novel reconstruction. This outline serves as a prompt to various baseline systems, as shown in Table  2.
This setting tests the upper bound of compression and allows us to observe distortion when expanding from extremely limited context.
@@ 3.2 Hierarchical Setup ($K=2$)
We next implement a two-stage outline generation pipeline inspired by rate-distortion theory. The original novel $X_{0}$ is first compressed into a global outline $X_{1}$, then each major section is further compressed into a section-level outline $X_{2}$. This structure enables multi-resolution control over narrative details.
Compression ratios $\alpha_{1}\in\{0.05,0.10\}$ and $\alpha_{2}\in\{0.01,0.05,0.10,0.20\}$ determine the abstraction levels, as shown in Table 4..
@@
Outlines is structured using JSON templates paragraph summaries, plain text will be tested in controlled ablations.
@@ 4.1 Significance Testing
We conducted pairwise $t$-tests to evaluate the effect of compression settings on novel reconstruction distortion. Results indicate that most configuration pairs exhibit statistically significant differences ($p<0.05$), suggesting that compression parameters substantially influence distortion, especially character similarity. Notably, K2-* outperforms all other configurations ($p<0.001$), whereas K2-3 underperforms compared to K2-4 and K2-5 in terms of character similarity. A few comparisons (e.g., K2-1 vs. K2-2, $p=0.60$) show no significant difference, indicating robustness within certain parameter ranges. B shows that translated text is very similar to original text.
Ablation Study.
For each variable (e.g., $\alpha_{1}$, $\alpha_{2}$, prompt format, temperature), we perform pairwise controlled comparisons while holding others fixed. Work in Progress.
@@
We present: (i) The correlation between the compression ratio product $R$ and the average similarity score, computed as $1-D(\hat{X},X_{0})$, as shown in Figure 3 in Appendix. (ii)Grouped statistics (mean ± std) of similarity and structure difference metrics under different compression settings, as shown in Table 4. (iii)Pairwise Significance Test, as shown in Figure 2 in Appendix.

### v2

# Measuring Information Distortion in Hierarchical Ultra long Novel Reconstruction: The Optimal Expansion Ratio

@@ Abstract
A two stage novel generation framework (outline -> section outline -> manuscript) is widely used in long novel generation,(e.g., DOME, Plan&Write, Long Writer), but study of such framework in ultra long novel(>1M words) reconstruction is little.
Building on recent text compression methods (LLMZip, LLM2Vec), we conduct an information-theoretic analysis to quantify semantic distortion under different compression–expansion ratios. We examine how outline length affects information preservation.
Experiments on ultra long novels show that the optimal compression-expansion ratio significantly reduce semantic distortion compare to other non-optimal compression-expansion ratio.
Measuring Information Distortion in Hierarchical Ultra long Novel Reconstruction:
@@ 1 Introduction
We observe a fundamental phenomenon in large-scale novel generation (Li et al., 2022; Su et al., 2022): when a million-word novel is summarized into a 1,000-word outline by an LLM (e.g., ChatGPT) and then expanded, the result shows substantial distortion, detail loss, and semantic divergence. In contrast, compressing a 100k-word novel into a 10k-word outline and expanding it preserves meaning more faithfully.
Studies such as LongWriter agree that generating 100k+ word novels typically follows an outline-to-novel workflow. The quality depends on two prompt-side variables: (i) outline length, and (ii) detail density (e.g., characters, plots, scenes). A trade-off emerges: fewer tokens and sparse detail reduce quality, while rich, lengthy prompts shift the burden to the human author.
Our ultimate goal is to generate detail-rich ultra-long novels from minimal outlines. However, since "detail" itself is hard to define and evaluate directly, we instead study the detail loss in reconstruction of ultra-long novels under varying compression ratios using large language models, as a proxy for generation quality.
@@ 1.1 Motivation
Ultra-long novels (1M+ words) are highly popular on WuxiaWorld, Fanqie, and Qidian, making automated generation a key goal. Since ChatGPT, models like LLaMA (Grattafiori et al., 2024), DeepSeek (DeepSeek-AI et al., 2025), Qwen (Bai et al., 2023), Gemini (Team et al., 2024), and GPT-4o (OpenAI et al., 2024) have advanced long-context processing.
However, despite 1M-token input capacity, output limits (e.g., 16k) make faithful reconstruction difficult (Mikhaylovskiy, 2023). Although many prior studies have addressed novels up to 100k words, empirical evidence is lacking to show that methods effective at this scale naturally extend to generating novels of 1 million words. Challenges at the million-word scale—maintaining coherence, thematic consistency, and character development—are qualitatively and quantitatively different. Inspired by the encoder–decoder paradigm, we adopt a reconstruction-based framework as a surrogate objective to study and improve ultra-long text generation.
@@ 1.2 Contribution
  • We establish a quantitative relationship between outline compression-expansion ratios and distortion (detail loss) through experiments. $R=0.01$ is the optimal compression-expansion ratio under our configuration and experiment.
  • We propose and empirically verify that mixed two-stage hierarchical outline decreases distortion in generated novels compared to traditional one-stage and two-stage approaches.
@@ 1.3 Related Work
Long-text generation introduces challenges in coherence, planning, and discourse modeling (Kumar et al., 2024; Que et al., 2024; Wu et al., 2025). Techniques include retrieval-augmented generation (Lewis et al., 2021), self-refinement (Madaan et al., 2023), and iterative lengthening (Quan et al., 2024). Novel generation demands plot and character consistency over long spans (Xie et al., 2023; Yao et al., 2019; Guan et al., 2021). Hierarchical planning (Wang et al., 2023; Yang et al., 2023; Wang et al., 2024), memory-based agents (Bai et al., 2024), and reinforcement-based control (Doshi and Hauser, 2024; Chhun, ) have been explored, but ultra-long contexts remain difficult.
@@
Some work treats LLMs as adaptive entropy coders (Shin, ), e.g., LLMZip (Valmeekam et al., 2023) and DeepZip (Goyal et al., 2018), achieving near-Shannon compression via token probability coding. We instead explore lossy but semantically faithful compression: can an LLM expand a sparse outline into a coherent million-word novel? We analyze the acceptable range of outline entropy $H_{\text{outline}}$ that balances compression with reconstructive fidelity.
@@
We assess both (i) distortion from the original novel and (ii) the quality of generated output. Distortion metrics include traditional algorithms (Cosine, Hamming, Jaccard, Levenshtein, TF-IDF) (Gahman and Elangovan, 2023a; Gomaa and Fahmy, 2013; Singh and Singh, 2020; Wang and Dong, 2020; Gahman and Elangovan, 2023b) and pre-trained models (BERTScore, BLEU, ROUGE, METEOR) (Guan et al., 2022; Xiao et al., 2024; Yu et al., 2024; Yuan et al., 2023). Novel quality is evaluated via aesthetic score, coherence, creativity, consistency, and fluency (Ichmoukhamedov et al., 2024; Venkatraman et al., 2024), with recent studies proposing LLMs themselves as evaluators (Gu et al., 2025).
@@ 2 Methodology
We implement three pipelines, as shown in Figure 1: (i)Single-Stage: generated from the full novel, then expanded directly to explore distortion under extreme compression. (ii)Hierarchical Two-Stage: The compression first reduces the original novel $X_{0}$ to a high-level outline $X_{1}$ with rate $\alpha_{1}$, and then further compresses it to a global summary $X_{2}$ with rate $\alpha_{2}$. The expansion reconstructs $\hat{X}_{1}$ from $X_{2}$, and finally recovers $\hat{X}_{0}$ from $\hat{X}_{1}$.
(iii)Mixed Two-Stage: The compression is the same as above. However, the expansion skips the intermediate outline and directly reconstructs $\hat{X}_{0}$ from $X_{2}$, under varying compression ratios $R$.
We can derive the compression ratio to be $R=\alpha_{1}\times\alpha_{2}$.
Using traditional methods like LZ77, lossless compression achieves $R\approx 0.5$ (Amir et al., 2024), which serves as a lower bound.
We define distortion as a composite of Cosine distance, BERTScore, and Euclidean distance of detail counts, detailed in our evaluation metrics.
@@
K2-6 0.015 0.668 $\pm$ 0.073 0.193 $\pm$ 0.041 0.667 $\pm$ 0.184 0.535 $\pm$ 0.167 0.628 $\pm$ 0.119 8.76 $\pm$ 6.56 4.98 $\pm$ 3.94 5.03 $\pm$ 6.07
K2-7 0.020 0.668 $\pm$ 0.07 0.192 $\pm$ 0.044 0.71 $\pm$ 0.166 0.583 $\pm$ 0.158 0.654 $\pm$ 0.110 9.53 $\pm$ 7.58 5.09 $\pm$ 4.30 5.42 $\pm$ 7.00
K2-8 0.015 0.674 $\pm$ 0.075 0.198 $\pm$ 0.043 0.67 $\pm$ 0.181 0.531 $\pm$ 0.168 0.634 $\pm$ 0.116 8.47 $\pm$ 6.59 4.63 $\pm$ 3.81 4.49 $\pm$ 4.11
K2-9 0.020 0.679 $\pm$ 0.073 0.197 $\pm$ 0.045 0.663 $\pm$ 0.193 0.543 $\pm$ 0.202 0.633 $\pm$ 0.125 9.41 $\pm$ 6.23 5.17 $\pm$ 4.15 5.31 $\pm$ 5.03
Table 2: Grouped statistics (mean $\pm$ std) for similarity and structural difference metrics, excluding the baseline (B), to identify best-performing configurations. The compression ratio is computed as $R=\alpha_{1}\times\alpha_{2}$. Bolded values indicate the best average performance for each metric among the tested configurations under $R\leq 0.01$, but do not imply statistical significance. Formal significance testing results are presented in a separate table. Distortion is computed as one minus similarity.
2.2 Dataset Configuration
We sample 40 public-domain Chinese novels (10 each from Fantasy, Urban, Romance, and Historical genres) from dataset (zxbsmk, 2023). Each 1M-word novel is split into  5,000-word chapters for reliable processing.
Compression/expansion prompts are detailed in Appendix CD. Translation experiments confirm language versions have minimal impact.
For evaluation (Appendix J), we generate 200 section and one global outline per novel. Eight chapters per novel are reconstructed for comparison.
2.3 Controlled Variables.
Our configuration space is defined by five independent factors: (i) the novel genre selected for evaluation; (ii) the compression ratio applied at each outline level (global and section); (iii) the LLM model, which is the backbone model.
2.4 Evaluation Metrics
To quantify reconstruction distortion, we define a three distance function $D(X,\hat{X})$ as:
$D_{\text{trad}}(X,\hat{X})$ combines one minus the cosine similarity score and one minus BERTScore between embeddings of original and reconstructed novels;
$D_{\text{llm}}(X,\hat{X})$ is the average GPT-4o judgment score across semantic, plot, character, background, and style similarity;
$D_{\text{struct}}(X,\hat{X})$ measures absolute differences in counts of unique characters, scenes, and items.
@@ 3 Experiment
We evaluate hierarchical compression–expansion strategies for ultra-long novel generation using Gemini 2.0 Flash (temperature = 0.3), focusing on: (1) How do different compression ratios affect semantic fidelity? (2) Does a two-stage outline hierarchy ($K=2$) outperform a single-stage ($K=1$)?
Gemini is chosen for its large context window and throughput efficiency. Each novel involves near 3M input and 500K output tokens.
@@ 3.1 Baseline Setup ($K=1$)
We test a single-stage pipeline using a 10,000-word outline ($R=\alpha_{1}\approx 0.01$) generated from the full novel, then expanded directly to explore distortion under extreme compression.
We compare four methods: (A) Human-written original, (B) Translate → Reconstruct, (C) Direct compression–expansion, and (D) LongWriter.
@@ 3.2 Hierarchical Setup ($K=2$)
In the two-stage pipeline, the novel $X_{0}$ is compressed into a section-level outline $X_{1}$, then further into a global outline $X_{2}$, enabling multi-resolution control.
Compression ratios $\alpha_{1}\in\{0.05,0.10\}$, $\alpha_{2}\in\{0.01,0.05,0.10,0.20,0.30,0.40\}$ determine abstraction levels (Table 3).
@@
Outlines is structured using JSON templates paragraph summaries.
@@ 4.1 Significance Testing
Pairwise $t$-tests show that most compression configurations differ significantly in distortion ($p<0.05$), especially in character similarity. Notably, K2-* outperforms all settings with $R\leq 0.01$ ($p<0.001$), while K2-3 underperforms compared to K2-4 and K2-5. Some comparisons (e.g., K2-1 vs. K2-2, $p=0.60$) show no significant difference, suggesting robustness in certain ranges. B confirms that translated text remains semantically close to the original.
Distortion decreases with higher compression ratio, but gains from $R=0.001$ to 0.010.010.01 are larger than those from 0.010.010.01 to 0.020.020.02.
@@
Appendix figures show: (i) correlation between $R$ and similarity (Figure 3); (ii) grouped statistics (Table 3); (iii) significance heatmap (Figure 2).